# Machine-Learning Characterization of Intermittency in Plasma Turbulence: Single and Double Sheet Structures

### v2

@@ Abstract
software: JAX (Bradbury et al., 2018), numba (Lam et al., 2015), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), aweSOM (Ha et al., 2024, submitted)
@@ 1 Introduction
We employ a theory-agnostic way of characterizing the intermittency in magnetically dominated plasma turbulence. We study the nature of the intermittency in three-dimensional (3D) plasma turbulence using a combination of machine learning (ML) and computer vision techniques. Namely, we identify structures consistent with current sheets via a non-linear $N$-dimensional correlation with a CPU/GPU-accelerated implementation of self organizing maps (SOM; Kohonen, 1990). Then, we combine multiple realizations of the SOM model via a statistical ensemble learning framework to obtain statistically more reliable clustering solutions. Intermittent structures in plasma turbulence have previously been identified and their statistics analyzed with the structure functions (Davis et al., 2024), thresholding of current densities (Zhdankin et al., 2013; Kadowaki et al., 2018), Gaussian-fitting techniques (Greco et al., 2009; Chhiber et al., 2020), and, recently, ML-based segmentation (Serrano et al., 2024). Previous attempts at leveraging ML methods to detect current sheets include the $K$-means clustering and DBSCAN methods (Sisti et al., 2021), convolutional neural networks (Hu et al., 2020), and SOM (Bussov & Nättilä, 2021; Köhne et al., 2023; Edmond et al., 2024). These studies offer a promising outlook on the usability of novel computational methods to identify, in particular, current sheets in 2D simulations and in magnetospheric observations. Our study is the first of its kind that applies such technique to 3D simulations of highly-magnetized, relativistic plasma turbulence.
@@ 2 Simulation
The technical details of the simulation are described in Nättilä & Beloborodov (2021). Briefly, the simulation is a cubic box of size $L_{\rm sim}=426\;c/\omega_{\rm p}$, covering 1280 grid cells per dimension. The freely evolving system is initially in an unperturbed equilibrium state with neutral $e^{\pm}$-pair plasma, which is magnetized with a uniform external field $\mathbf{B_{0}}=B_{0}\hat{\mathbf{z}}$. We focus on the strongly magnetized domain where the magnetization parameter $\sigma\equiv B_{0}^{2}/4\pi nm_{e}c^{2}\gtrsim 1$ (where $n$ is the plasma number density, $m_{e}$ is the electron rest mass, and $c$ is the speed of light). We set $\sigma=10$. We consider the simplest pair plasma composition to minimize the role of the kinetic effects, with 54 particles per $c^{3}/\omega_{\rm p}^{3}$ per species (two particles per cell per species). The initial equilibrium is disturbed by exciting large-scale ($=l_{0}=L_{\rm sim}/4$) magnetic perturbations, $\delta\mathbf{B}\perp\mathbf{B_{0}}$ with a large amplitude $\delta B/B_{0}\sim 1$. The resulting plasma bulk motions are trans-relativistic, since $|\mathbf{u}|\approx v_{A}\approx c$, where $v_{A}\equiv c\sqrt{\sigma/(\sigma+1)}$ is the Alfvén velocity. We follow the simulations until $t=20\,l_{0}/c$ (where $l_{0}/c$ is the eddy-turnover time of the energy-carrying scale). We store data snapshots (composed of $\mathbf{E}$, $\mathbf{B}$, and $\mathbf{J}$ fields) on a cadence of $\Delta t\approx 1\,l_{0}/c$. Numerically, each snapshot we analyze is a down-sampled rectangular cube composed of $L^{3}=128^{3}\approx 2\times 10^{6}$ data points. Figure 1 shows a volume rendering of the projected current density along the $\mathbf{B}$ field, $j_{\parallel}\equiv\mathbf{J}\cdot\mathbf{B}/(Bn_{0}ec)$, together with a $xy$-slice plot of the same feature, over-plotted by streamlines of the in-plane magnetic field, $B_{\perp}$. We present the following results based on one representative snapshot of the simulation, taken at $t=5\;l_{0}/c$, which exhibits a magnetic power spectrum $\propto k_{\perp}^{\alpha}$ with a slope of $\alpha\approx-5/3$, where $k_{\perp}$ is the wavenumber perpendicular to the mean magnetic field.
Figure 1: Top: volume rendering of the projection of the current density onto the $\mathbf{B}$ field, $j_{\parallel}$. Middle: $xy$-slice of $j_{\parallel}$ at $z=90\,c/\omega_{\rm p}$. Streamlines of the in-plane magnetic field $b_{\perp}$ are in the foreground. Bottom: fiducial SOM segmentation, color-coded by the cluster identified via aweSOM. From observation and the radar charts in Figure 2, green and orange correspond to regions in between double current sheets, purple to current sheets aligned with $\mathbf{B}$, and blue to current sheets anti-aligned with $\mathbf{B}$. Blue contours are regions where $j_{\parallel}>2\,j_{\parallel,\rm rms}$, and red contours are regions where $j_{\parallel}<-2\,j_{\parallel,\rm rms}$.
@@ 3 Clustering Analysis
We analyze the resulting large, multidimensional dataset with a ML technique called Self-Organizing Maps (SOM; Kohonen, 1990). SOM is a sophisticated classification method that can capture multidimensional correlations in the input data by exposing the topology of the (possibly nonlinear) manifolds being analyzed. We employ the SOM algorithm specifically because it is a powerful, unsupervised method that does not require prior knowledge about the number of clusters (i.e., physical structures). Furthermore, the resulting clustering of nodes in a SOM is easily interpretable. We further enhance the robustness of the model by combining multiple SOM realizations through a statistically combined ensemble method (SCE; Bussov & Nättilä, 2021, see also Appendix C). SCE segmentation provides a statistically significant result by stacking SOM clusters of similar spatial distributions. We provide a numerically fast implementation of the SOM and SCE methods as an open-source Python package called aweSOM [^1] https://github.com/tvh0021/aweSOM/ (Ha et al., 2024, submitted). Specifically, the SOM implementation in aweSOM is an optimized and parallelized version of the R package POPSOM (Hamel, 2019), providing a marked improvement in training time, and vastly superior projection (the mapping of cluster label from lattice space to real space) time, on the order of $\sim 20-50$ times.
@@
Our fiducial SOM has a size of $64\times 37$ nodes in the lattice space. The map is trained with an initial learning rate $\alpha_{0}=0.1$ for $N=L^{3}\approx 2\times 10^{6}$ steps. We have found the reported map size to be the minimum viable option for adequately resolving the multidimensional feature space. We use a map with an aspect ratio of $H=37/64\approx 0.6$. The aspect ratio is selected to be close to the ratio of the first and second largest eigenvalues of the input data’s principal components. A detailed discussion of the technical parameters is given in Appendix A. During our extensive tests (see Appendix B), we found that the larger map direction tends to orient preferably along a latent space axis that strongly correlates with the system’s dominant principal component; physically, this axis is a combination of various components of $\mathbf{E}$, $\mathbf{B}$, and $\mathbf{J}$. The shorter map direction, on the other hand, tends to separate the positive and negative data values; physically, this can be interpreted as separating the directions aligned and anti-aligned with $\mathbf{B_{0}}$.
@@ 4.1 Fiducial SOM realization
Altogether, clusters 0 and 4 are single current sheets in isolation, but when detected adjacent to clusters 1 and 2, they are components of double current sheets. Taking a $xy$-slice as seen in Figure 1, we point to the structure at $x=90\,c/\omega_{\rm p}$ and $y=400\,c/\omega_{\rm p}$ as an example of a single current sheet, while at $x=350\,c/\omega_{\rm p}$ and $y=150\,c/\omega_{\rm p}$ a double current sheet is found. The volume-filling fraction, $n_{\rm fill}$, of these clusters are (for cluster 0, 1, $\ldots$): $5.4\%$, $1.7\%$, $1.1\%$, $2.5\%$, $3.3\%$, $1.5\%$, $84.5\%$. The regions with sheet-like structure contains a volume-filling fraction of $n_{\mathrm{fill,\,single+double}}\approx 11.5\%$ of the domain, which is in line with similar studies of strongly magnetized plasma turbulence (e.g., TenBarge & Howes, 2013; Vega et al., 2023). An additional statistical breakdown of each cluster is examined in Appendix E.
@@ 4.2 Separating single and double current sheets
Figure 3: Similar to the bottom panel of Figure 1, but with adjacent clusters 0, 1, 2, and 4 merged into cluster 1. Cluster 1 denotes double current sheets, cluster 4 and 0 denotes single current sheets that align with B and anti-align with B, respectively.
From the fiducial snapshot shown in Figure 1, we compute the volume-filling fraction of the single and double current sheets separately. Since cluster 1 and 2 only occupy the spaces in between peaks of the double current sheets, we merge all continuous double sheets clumps with their adjacent single current sheets (cluster 0 and 4) into complete double current sheets. The remaining clumps from cluster 0 and 4 that are not merged are true single current sheets. Lastly, clumps identified as cluster 1 or 2 but are not found to be adjacent to both a $-j_{\rm peak}$ and a $+j_{\rm peak}$ are considered false detections, and are merged into the background cluster.
Figure 3 shows an $xy$-slice of the clustering result after merging, keeping the same color notations as in Figures 1 and 2. Here, cluster 1 (orange) represents double current sheets; cluster 4 (purple) and cluster 0 (blue) are single current sheets that are aligned and anti-aligned with B, respectively; cluster 6 (white) is the background plasma; cluster 3 (red) and 5 (brown) are still locations of strong alignment between E and B. Altogether, $n_{\rm fill,\,double}\approx 9.1\%$ and $n_{\rm fill,\,single}\approx 1.5\%$. When the system is allowed to reach steady state without continuous energy injection, we find that most “current sheet” structures found are double current sheets. On the other hand, in continuously-driven turbulent environments, the relative fraction of single current sheets is more comparable to that of double current sheets ($n_{\rm fill,\,double}\approx 0.91\%$ and $n_{\rm fill,\,single}\approx 0.46\%$ in our analysis; see Appendix F).
@@ 4.3 SCE result
Figure 4: SCE result obtained by combining the output of 36 SOM realizations. The left panel shows a $xy$-slice plot of $\Sigma G_{\rm sum}$ (signal strength) at $z=90\,c/\omega_{\rm p}$. The right panels show SCE clusters, obtained via making thresholds in the signal strength at 0.25, 0.5, and 0.8, respectively. Red and blue contours are the same as in Figure 1, and denotes the regions where $|j_{\parallel}|>2\,j_{\parallel,\rm rms}$. Nominally, cluster 0 (pink) are double current sheets, cluster 1 (gray) are single current sheets aligned with $\mathbf{B}$, cluster 2 (olive) are single current sheets anti-aligned with $\mathbf{B}$, and cluster 3 (teal) are background plasma.
@@
We quantitatively verify the robustness of these clustering results by conducting an SCE analysis on 36 different SOM realizations. The SCE clusters are derived by combining the strongest common features across the SOM realizations—features that are more universally present are indicated by stronger signals. Figure 4 shows a $xy$-slice plot of the cumulative SCE $G_{\rm sum}$ values obtained via cluster-to-cluster comparison of all clusters $C$ in all aweSOM realizations $R$ (see Appendix C for a more detailed definition of $G_{\rm sum}$). The prominent clusters reveal four main structures: individual current sheets in olive and gray, double current sheets in pink, and the background plasma in teal.
@@ 4.4 Intermittency and current sheets
Interestingly, we find that some of the identified clusters are associated with spatio-temporal fluctuations in the simulation domain (clusters 0, 1, 2, and 4 in our fiducial analysis visualized in Figure 1 and 2 and cluster 0, 1, and 2 in Figure 4). Physically, these clusters coincide with localized patches of intense electric current: current sheets. Because of this, we associate these clusters with MHD intermittency. However, we always find two differing sheet-like structures in the flow. This is in stark contrast with the previous analyses (e.g., Chernoglazov et al., 2021; Fielding et al., 2023; Davis et al., 2024) that have assumed all current sheets to be of the same origin.
Figure 5 shows an example of the two types of current sheets we regularly identify with the SOM analysis. The inset plot in the top panels shows a $xy$-slice plot of $j_{\parallel}$, overplotted with a line segment $\overline{AB}$, which is approximately normal to the sheet(s). We denote the direction along $\overline{AB}$ as $\mathbf{n}$, and the direction along $\mathbf{\hat{z}}\times\mathbf{n}$ as $\mathbf{p_{1}}$. Then, we compute the line profile of $j_{\parallel}$, $B_{p_{1}}$, $B_{n}$, and $\mathbf{J}\cdot\mathbf{E}$ along $\overline{AB}$, and plot the profiles on the top panels.

### v3

# Machine-Learning Characterization of Intermittency in Relativistic Pair Plasma Turbulence: Single and Double Sheet Structures

@@ Abstract
journal: ApJLsoftware: JAX (Bradbury et al., 2018), numba (Lam et al., 2015), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), aweSOM(Ha et al., 2025)
@@ 1 Introduction
We employ a theory-agnostic way of characterizing the intermittency in magnetically dominated plasma turbulence. We study the nature of the intermittency in three-dimensional (3D) plasma turbulence using a combination of machine learning (ML) and computer vision techniques. Namely, we identify structures consistent with current sheets via a non-linear $N$-dimensional correlation with a CPU/GPU-accelerated implementation of self organizing maps (SOM; Kohonen, 1990). Then, we combine multiple realizations of the SOM model via a statistical ensemble learning framework to obtain statistically more reliable clustering solutions. Intermittent structures in plasma turbulence have previously been identified and their statistics analyzed with the structure functions (Davis et al., 2024), thresholding of current densities (Zhdankin et al., 2013; Kadowaki et al., 2018), Gaussian-fitting techniques (Greco et al., 2009; Chhiber et al., 2020), and, recently, ML-based segmentation (Serrano et al., 2025). Previous attempts at leveraging ML methods to detect current sheets include the $K$-means clustering and DBSCAN methods (Sisti et al., 2021), convolutional neural networks (Hu et al., 2020), and SOM (Bussov & Nättilä, 2021; Köhne et al., 2023; Edmond et al., 2024). These studies offer a promising outlook on the usability of novel computational methods to identify, in particular, current sheets in 2D simulations and in magnetospheric observations. Our study is the first of its kind that applies such technique to 3D simulations of highly-magnetized, relativistic plasma turbulence.
@@ 2 Simulation
Figure 1: Panel a): volume rendering of the projection of the current density onto the $\mathbf{B}$ field, $j_{\parallel}$. Panel b): $xy$-slice of $j_{\parallel}$ at $z=90\,c/\omega_{\rm p}$. Streamlines of the in-plane magnetic field $b_{\perp}$ are in the foreground. Panel c): fiducial SOM segmentation, color-coded by the cluster identified via aweSOM. From observation and the radar charts in Figure 2, green and orange correspond to regions in between double current sheets, purple to current sheets aligned with $\mathbf{B}$, and blue to current sheets anti-aligned with $\mathbf{B}$. Blue contours are regions where $j_{\parallel}>2\,j_{\parallel,\rm rms}$, and red contours are regions where $j_{\parallel}<-2\,j_{\parallel,\rm rms}$.
@@
The technical details of the simulation are described in Nättilä & Beloborodov (2021). Briefly, the simulation is a cubic box of size $L_{\rm sim}=426\;c/\omega_{\rm p}$, covering 1280 grid cells per dimension. The freely evolving system is initially in an unperturbed equilibrium state with neutral $e^{\pm}$-pair plasma, which is magnetized with a uniform external field $\mathbf{B_{0}}=B_{0}\hat{\mathbf{z}}$. We focus on the strongly magnetized domain where the magnetization parameter $\sigma\equiv B_{0}^{2}/4\pi nm_{e}c^{2}\gtrsim 1$ (where $n$ is the plasma number density, $m_{e}$ is the electron rest mass, and $c$ is the speed of light). We set $\sigma=10$. We consider the simplest pair plasma composition to minimize the role of the kinetic effects, with 54 particles per $c^{3}/\omega_{\rm p}^{3}$ per species (two particles per cell per species). The initial equilibrium is disturbed by exciting large-scale ($=l_{0}=L_{\rm sim}/4$) magnetic perturbations, $\delta\mathbf{B}\perp\mathbf{B_{0}}$ with a large amplitude $\delta B/B_{0}\sim 1$. The resulting plasma bulk motions are trans-relativistic, since $|\mathbf{u}|\approx v_{A}\approx c$, where $v_{A}\equiv c\sqrt{\sigma/(\sigma+1)}$ is the Alfvén velocity. We follow the simulations until $t=20\,l_{0}/c$ (where $l_{0}/c$ is the eddy-turnover time of the energy-carrying scale). We store data snapshots (composed of $\mathbf{E}$, $\mathbf{B}$, and $\mathbf{J}$ fields) on a cadence of $\Delta t\approx 1\,l_{0}/c$. Numerically, each snapshot we analyze is a down-sampled rectangular cube composed of $L^{3}=128^{3}\approx 2\times 10^{6}$ data points. Figure 1a shows a volume rendering of the projected current density along the $\mathbf{B}$ field, $j_{\parallel}\equiv\mathbf{J}\cdot\mathbf{B}/(Bn_{0}ec)$, together with a $xy$-slice plot of the same feature, over-plotted by streamlines of the in-plane magnetic field, $B_{\perp}$ in Figure 1b. We present the following results based on one representative snapshot of the simulation, taken at $t=5\;l_{0}/c$, which exhibits a magnetic power spectrum $\propto k_{\perp}^{\alpha}$ with a slope of $\alpha\approx-5/3$, where $k_{\perp}$ is the wavenumber perpendicular to the mean magnetic field.
@@ 3 Clustering Analysis
We analyze the resulting large, multidimensional dataset with a ML technique called Self-Organizing Maps (SOM; Kohonen, 1990). SOM is a sophisticated classification method that can capture multidimensional correlations in the input data by exposing the topology of the (possibly nonlinear) manifolds being analyzed. We employ the SOM algorithm specifically because it is a powerful, unsupervised method that does not require prior knowledge about the number of clusters (i.e., physical structures). Furthermore, the resulting clustering of nodes in a SOM is easily interpretable. We further enhance the robustness of the model by combining multiple SOM realizations through a statistically combined ensemble method (SCE; Bussov & Nättilä, 2021, see also Appendix C). SCE segmentation provides a statistically significant result by stacking SOM clusters of similar spatial distributions. We provide a numerically fast implementation of the SOM and SCE methods as an open-source Python package called aweSOM [^1] https://github.com/tvh0021/aweSOM/ (Ha et al., 2025). Specifically, the SOM implementation in aweSOM is an optimized and parallelized version of the R package POPSOM (Hamel, 2019), providing a marked improvement in training time, and vastly superior projection (the mapping of cluster label from lattice space to real space) time, on the order of $\sim 20-50$ times faster.
@@
Our fiducial SOM has a size of $64\times 37$ nodes in the lattice space. The map is trained with an initial learning rate $\alpha_{0}=0.1$ for $N=L^{3}\approx 2\times 10^{6}$ steps. We have found the reported map size to be the minimum viable option for adequately resolving the multidimensional feature space. We use a map with an aspect ratio of $H=37/64\approx 0.6$. The aspect ratio is selected to be close to the ratio of the first and second largest eigenvalues of the input data’s principal components. A detailed discussion of the technical parameters is given in Appendix A. During our extensive tests (see Appendix B and E, we found that the SOM lattice encompasses a surface in the 4D feature space, which is not easily translated to a physical connection. For example, Figure 7 shows projections of the lattice on pairs of features. The nodes along $Y=0$ (the black line going from green to purple) span a wide range of values in $j_{\rm a}$, but does not show a similar behavior in any other feature. At the same time, the nodes along $Y=36$ (the black line going from blue to brown) span both extremes in $e_{\parallel}$. The shorter lattice direction ($Y$), on the other hand, tends to separate the positive and negative data values; physically, this can be interpreted as separating the directions aligned and anti-aligned with $\mathbf{B_{0}}$.
@@ 4.1 Fiducial SOM realization
Altogether, clusters 0 and 4 are single current sheets in isolation, but when detected adjacent to clusters 1 and 2, they are components of double current sheets. Taking a $xy$-slice as seen in Figure 1b, we point to the structure at $x=90\,c/\omega_{\rm p}$ and $y=400\,c/\omega_{\rm p}$ as an example of a single current sheet, while at $x=350\,c/\omega_{\rm p}$ and $y=150\,c/\omega_{\rm p}$ a double current sheet is found. The volume-filling fraction, $n_{\rm fill}$, of these clusters are (for cluster 0, 1, $\ldots$): $5.4\%$, $1.7\%$, $1.1\%$, $2.5\%$, $3.3\%$, $1.5\%$, $84.5\%$. The regions with sheet-like structure contains a volume-filling fraction of $n_{\mathrm{fill,\,single+double}}\approx 11.5\%$ of the domain, which is in line with similar studies of strongly magnetized plasma turbulence (e.g., TenBarge & Howes, 2013; Vega et al., 2023). An additional statistical breakdown of each cluster is examined in Appendix E.
@@ 4.2 Separating single and double current sheets
Figure 3: Similar to panel c) of Figure 1, but with adjacent clusters 0, 1, 2, and 4 merged into cluster 1. Cluster 1 denotes double current sheets, cluster 4 and 0 denotes single current sheets that align with B and anti-align with B, respectively.
From the fiducial snapshot shown in Figure 1c, we compute the volume-filling fraction of the single and double current sheets separately. Since cluster 1 and 2 only occupy the spaces in between peaks of the double current sheets, we merge all continuous double sheets clumps with their adjacent single current sheets (cluster 0 and 4) into complete double current sheets. The remaining clumps from cluster 0 and 4 that are not merged are true single current sheets. Lastly, clumps identified as cluster 1 or 2 but are not found to be adjacent to both a $-j_{\rm peak}$ and a $+j_{\rm peak}$ are considered false detections, and are merged into the background cluster.
Figure 3 shows an $xy$-slice of the clustering result after merging, keeping the same color notations as in Figures 1c and 2. Here, cluster 1 (orange) represents double current sheets; cluster 4 (purple) and cluster 0 (blue) are single current sheets that are aligned and anti-aligned with B, respectively; cluster 6 (white) is the background plasma; cluster 3 (red) and 5 (brown) are still locations of strong alignment between E and B. Altogether, $n_{\rm fill,\,double}\approx 9.1\%$ and $n_{\rm fill,\,single}\approx 1.5\%$. When the system is allowed to reach steady state without continuous energy injection, we find that most “current sheet” structures found are double current sheets. On the other hand, in continuously-driven turbulent environments, the relative fraction of single current sheets is more comparable to that of double current sheets ($n_{\rm fill,\,double}\approx 0.91\%$ and $n_{\rm fill,\,single}\approx 0.46\%$ in our analysis; see Appendix F).
@@ 4.3 SCE result
Figure 4: SCE result obtained by combining the output of 36 SOM realizations. The left panel shows a $xy$-slice plot of $\Sigma G_{\rm sum}$ (signal strength) at $z=90\,c/\omega_{\rm p}$. The right panels show SCE clusters, obtained via making thresholds in the signal strength at 0.25, 0.5, and 0.8, respectively. Red and blue contours are the same as in Figure 1c, and denotes the regions where $|j_{\parallel}|>2\,j_{\parallel,\rm rms}$. Nominally, cluster 0 (pink) are double current sheets, cluster 1 (gray) are single current sheets aligned with $\mathbf{B}$, cluster 2 (olive) are single current sheets anti-aligned with $\mathbf{B}$, and cluster 3 (teal) are background plasma.
@@
We quantitatively verify the robustness of these clustering results by conducting an SCE analysis on 36 different SOM realizations. The SCE clusters are derived by combining the strongest common features across the SOM realizations—features that are more universally present are indicated by stronger signals. Figure 4 shows a $xy$-slice plot of the cumulative SCE $G_{\rm sum}$ values obtained via cluster-to-cluster comparison of all clusters $C$ in all aweSOM realizations $R$ (see Appendix C for a more detailed definition of $G_{\rm sum}$). The prominent clusters reveal four main structures: individual current sheets in olive and gray, double current sheets in pink, and the background plasma in teal. We caution that, since the SCE analysis is run on the original SOM clustering results without post-processing, many of the individual sheets identified by the SCE algorithm are not of single type, but instead parts of the double-sheet structures if detected adjacent to cluster 0 (pink). The more accurate statistical representation of the two types of intermittent structures is presented in Section 4.2.
@@ 4.4 Intermittency and current sheets
Interestingly, we find that some of the identified clusters are associated with spatio-temporal fluctuations in the simulation domain (clusters 0, 1, 2, and 4 in our fiducial analysis visualized in Figure 1c and 2 and cluster 0, 1, and 2 in Figure 4). Physically, these clusters coincide with localized patches of intense electric current: current sheets. Because of this, we associate these clusters with MHD intermittency. However, we always find two differing sheet-like structures in the flow. This is in stark contrast with the previous analyses (e.g., Chernoglazov et al., 2021; Fielding et al., 2023; Davis et al., 2024) that have assumed all current sheets to be of the same origin.
Figure 5 shows an example of the two types of current sheets we regularly identify with the SOM analysis. The inset plot in the top panels shows a $xy$-slice plot of $j_{\parallel}$, overplotted with a line segment $\overline{AB}$, which is approximately normal to the sheet(s). We denote the direction along $\overline{AB}$ as $\mathbf{n}$, and the direction along $\mathbf{\hat{z}}\times\mathbf{n}$ as $\mathbf{p_{1}}$. Then, we compute the line profile of $j_{\parallel}$, $B_{p_{1}}$, $B_{n}$, and $\mathbf{J}\cdot\mathbf{E}$ along $\overline{AB}$, and plot the profiles in panels a) and c).